OpenStudy (anonymous):

Explain, in complete sentences, whether or not the polynomial x^2 – 16x + 64 fits one of the special patterns. If so, which one? Part 2: Explain how the polynomial is factored.

OpenStudy (anonymous):

flvs?

OpenStudy (anonymous):

This is a perfect square form. to test for a perfect square, take square root of the leading coefficient, then square root of the constant term, which is x and 8 respectively. Next step is to test of it is a perfect square form. To do so, calculate the factor 2 (x) (8) = 16x that matches the second term. (Ignore the sign of the second for this step. Finally, write (x - 8)^2 since the second term is negative. Keep in mind that to be a perfect square term, the first and the last term have to in perfect square form and thus positive.

OpenStudy (anonymous):

The special pattern is (a - b)^2 where a = x and b = 8

OpenStudy (anonymous):

what are the factors.

OpenStudy (anonymous):

(x - 8)^2 or (x - 8)(x - 8). We also can use factor by grouping. x^2 - 16x - 64 = x^2 - 8x - 8x +16 = x(x - 8) - 8(x - 8) = (x - 8)(x - 8)